OpenStudy (anonymous):

write y-10=-(x-2) in standard form?

OpenStudy (anonymous):

Standard form looks like y=mx+b, where m is the slope and b is the y intercept. First you need to distribute the -1 to (x-2) and you will get y-10=-x+2. Then, you add the 10 to the right side, so the standard form is y=-x+12.